# How many qubits does a machine learning problem require?

### v1

@@ Abstract
For a machine learning paradigm to be generally applicable, it should have the property of universal approximation, that is, it should be able to approximate any target function to any desired degree of accuracy. In variational quantum machine learning, the class of functions that can be learned depend on both the data encoding scheme as well as the architecture of the optimizable part of the model. Here, we show that the property of universal approximation is constructively and efficiently realized by the recently proposed bit-bit encoding scheme. Further, we show that this construction allows us to calculate the number of qubits required to solve a learning problem on a dataset to a target accuracy, giving rise to the first resource estimation framework for variational quantum machine learning. We apply bit-bit encoding to a number of medium-sized datasets from OpenML and find that they require only $20$ qubits on average for encoding. Further, we extend the basic bit-bit encoding scheme to one that can handle batching very large datasets. As a demonstration, we apply this new scheme to the giga-scale transcriptomic Tahoe-100M dataset, concluding that the number of qubits required for encoding it lies beyond classical simulation capabilities. Remarkably, we find that the number of qubits does not necessarily increase with the number of features of a dataset, but may sometimes even decrease.
@@ 1 Introduction
The use of quantum computers for machine learning and artificial intelligence applications is a relatively new research area that could provide a means to model datasets which contain correlations that are hard to model classically. Quantum machine learning algorithms have been tested on near-term quantum hardware for various real-world datasets [1, 2, 3, 4, 5, 6, 7, 8, 9]. In particular, variational quantum machine learning, which uses parameterized unitaries that can be trained to model the correlations in a dataset analogous to classical neural networks [10], can be applied to a wide variety of problems. However, compared to other applications such as chemistry [11, 12, 13, 14] or cryptography [15], where algorithms with well-understood quantum advantage and extensively studied resource estimation are known, there are many more open questions around how to use quantum computers effectively for learning from data.
@@
Among the fundamental questions is one of expressivity - how does one design a variational quantum ansatz, which we refer to as a quantum model, that can approximate the underlying function defined by a given dataset. To answer this question meaningfully, we must note that real-world datasets seldom follow theoretically tractable rules. Therefore, the quantum model should have the property of universal approximation - that is, it should be able to approximate any function with any desired degree of accuracy. Further, the construction which realizes the universal approximation should not itself require exponentially scaling computation, either quantum or classical. If a quantum learning model does not have this property, it is likely not useful in real life. Even if such a model works in a limited setting, we cannot expect it to give good results in the limit of large and complex datasets. Indeed, in classical machine learning, universal approximation theorems [16, 17] provide the mathematical justification for using neural networks for large and complex datasets.
@@
[18] showed that the expressivity of a quantum model is constrained by the data-encoding method used. Specifically, a quantum model can be expanded as a partial Fourier series in the encoded data where the frequencies present in the series are set by linear combinations of the eigenvalues of the encoding operators. Following this, it is possible to increase the number of frequencies in the series by uploading the data multiple times, a technique known as data reuploading. The time for loading is then lower-bounded by the quantity of information to be loaded multiplied by the number of times it is loaded. With this method, even an image from a simple dataset like MNIST, which has 784 pixels, is expensive to load directly into a quantum state. Since this limit is fundamental, it is clear that some kind of compression is required before loading, for which techniques like principal component analysis are often used. However, it remains unclear how much the degree of compression and the number of data reuploadings affect the accuracy that can be reached by the model.
@@
Here we show that the bit-bit encoding technique [19], which was recently proposed for supervised learning, directly addresses the problem of achieving efficient compression with predictable accuracy. In this encoding, both the input and the output are encoded as binary strings following a dimensionality reduction procedure. These binary strings correspond to the basis states of a quantum computer at initialization and measurement. For a small number of qubits, there are collisions in which the same input corresponds to more than one output, but as the number of qubits increases, the encoding becomes more precise. We use the Solovay-Kitaev theorem [20] to prove that a variational quantum model with bit-bit encoding has the property of universal approximation. Further, the classical pre-processing required before data loading utilizes techniques such as principal component analysis which scale only polynomially in the dataset size.
@@
A very useful emergent property of the bit-bit encoding construction is that it is possible to efficiently calculate exactly the number of qubits required to encode a dataset to a desired degree of accuracy. This, then gives us the first framework for resource estimation in quantum machine learning. Using this framework, we can for the first time estimate the number of qubits required to solve a learning problem using a variational approach. In turn, this analysis, has implications for assessing the potential of quantum advantage for a dataset as we can calculate whether or not a learning problem fits into a number of qubits that can be simulated on a classical computer.
@@
We introduce $Q_{\text{dataset}}$, a resource metric that estimates the number of qubits required to model a dataset using bit-bit encoding. Starting with the MNIST dataset, we illustrate how $Q_{\text{dataset}}$ is derived by incrementally increasing qubit allocations until collisions on both the train and test datasets vanish, revealing a trade-off between maximizing theoretical training accuracy and achieving generalization. Extending this analysis to a variety of OpenML datasets, we find that typical medium-sized classical machine learning datasets require $20$ qubits for coverage on average, suggesting they are unlikely to be candidates for quantum advantage [21]. Comparisons across dimensionality reduction methods show that schemes producing independent or de-correlated features achieve lower $Q_{\text{dataset}}$ than non-linear or unprocessed data. Finally, scaling experiments on the large Tahoe-100M transcriptomic dataset [22] reveal examples where quantum models require a number of qubits well beyond classical simulation capabilities, which peak at about 50 qubits [21], indicating that this is a good arena to look for quantum advantage.
@@
We note that we only consider classical datasets in this work. Although there has been progress in quantum machine learning for quantum data [23, 24], its practical applicability remains limited. In contrast, quantum machine learning applied to classical data offers a far broader range of potential applications..
@@
The paper is organized as follows. We begin with an overview of bit-bit encoding in Section 2, followed by an analysis of expressive power and universal approximation in commonly used encoding schemes in Section 3. Section 4 discusses several techniques that extend the original data encoding techniques proposed in paper [19], including techniques that can be used with large datasets. Section 4 also presents the number of qubits required for common datasets from OpenML. Section 5 demonstrates training a classifier on a bit-bit encoded dataset. We conclude with a discussion in Section 6.
2 Overview of Bit-Bit Encoding
@@
Bit-bit encoding [19] encodes datasets with numeric features and discrete class labels into bitstrings, which allows the dataset to be read into a Hillbert space of a given number of qubits for classification. Consider a supervised learning problem in which we are given a dataset with consists of a feature matrix $X\in\mathbb{R}^{s\times n}$ and class label vector $y\in\{0,1,\ldots,c\}^{s}$, where $s$ is the number of samples, $n$ is the number of features, and $c$ is the number of class labels. For a classification problem, the learning task consists of finding an efficiently computable function that maps $X$ to $y$. Typically, the function is many-to-one, that is, $c<|X|$.
@@
Bit-bit encoding converts the feature matrix $X$ into a discretized feature matrix $Z$, compactly encoded into $N_{x}$ qubits. The binary encoding algorithm works as follows. First, the dataset is split into train and test sets: $X_{\text{train}}$, $y_{\text{train}}$, $X_{\text{test}}$, and $y_{\text{test}}$. A dimensionality reduction model, principal component analysis (PCA) by default, is fit on $X_{\text{train}}$ and applied to $X_{\text{test}}$, reducing both sets to $D$ components. The number of bits assigned to each component is calculated with
@@
where $I$ is the mutual information and $N_{x}$ is the number of bits allocated to the discretized feature matrix. The dimension-reduced $X_{\text{train}}$ and $X_{\text{test}}$ are then min-max normalized so that each sample lies in $0,1]^{D}$. A copula transformation is fit on the reduced and normalized $X_{\text{train}}$ and applied to the reduced and normalized $X_{\text{test}}$, mapping each component to its empirical CDF so that its marginal distribution is approximately uniform within each class. Finally, to encode the reduced, normalized, and transformed $X_{\text{train}}$ and $X_{\text{test}}$, each data point $x_{sd}$ (the $s^{th}$ sample of the $d^{th}$ feature) is discretized into $b_{d}$ bits of precision with
@@
Discretizing the feature matrix also compresses it, potentially creating collisions at the input as a particular feature vector $z\in Z$ may occur more than once, each time corresponding to the same or different value of $y_{s}\in y$. The samples in the discretized feature matrix can then be considered as samples from a joint probability distribution $f(Z=z,y=y_{s})$ over correlated random variables $(Z,y)$. For the purpose of classification, the ‘correct’ output is then taken to be the mapping which occurs more frequently, that is
@@
The quantum learning problem can now be cast in terms of an operator $U_{*}$ that has the following action on $N_{q}=N_{x}+N_{y}$ qubits:
@@
find the value of $\vec{\theta}$ which minimizes the loss function
$\bar{L}(\vec{\theta})=1-\sum_{z\in\mathcal{D}_{\vec{b}}}f(z)P_{C(z),z}(\vec{\theta}),$ (3)
where $f(z)$ is the frequency of occurrence of $z$ in the dataset. $0\leq\bar{L}\leq 1$, and has the simple interpretation of being the probability of the model giving the wrong answer when queried. If $U(\vec{\theta})=U_{*}$ and there is no noise, even one shot would suffice to classify the input data sample.
Note that the loss function only includes the action of $U$ on values of $z$ in the dataset. The action of $U$ on values of $z$ not in the dataset will determine its generalization behavior. In that case, a quantum model that is perfectly trained on the dataset will output a probability distribution over the possible classes.
We distinguish the bit-bit encoding technique from a simple basis encoding as it can be used to represent datasets that do not originally have a binary representation, but are converted into binary representation through a process of dimensional reduction followed by mutual-information based discretization. Indeed, a basis encoding that utilized only a naive discretization of the dataset would require many more qubits. For instance, a dataset consisting of molecular fingerprint samples [25], which are naturally binary, would take thousands of qubits to encode.
@@ 3.1 Universality of bit-bit encoding
In this section, we show how the bit-bit encoding leads to a universal approximation theorem for classification. First, we prove that the classification operator $U_{*}$ defined in Eq. 1 is unitary.
@@
$C(z)=C(z^{\prime})\;\Longrightarrow\;\braket{g(z)|g(z^{\prime})}=\delta_{z,z^{\prime}}.$ (4)
@@
If $C(z)\neq C(z^{\prime})$ the first factor vanishes; if $C(z)=C(z^{\prime})$, the condition (4) gives $\braket{g(z)|g(z^{\prime})}=\delta_{z,z^{\prime}}$. Hence, in all cases,
$\bigl{\langle}C(z),g(z)\bigm{|}C(z^{\prime}),g(z^{\prime})\bigr{\rangle}=\delta_{z,z^{\prime}}.$ (7)
@@
is an orthonormal family in $\mathcal{H}$. Because $|\{0,1\}^{N_{x}}|=2^{N_{x}}$ and $\dim\mathcal{H}_{\mathrm{in}}=2^{N_{x}}$, we have $\dim\mathrm{span}(\mathcal{B}_{\mathrm{out}})=2^{N_{x}}$. Thus $V$ in (5) is an isometry from $\mathcal{H}_{\mathrm{in}}$ onto
@@
$\|U_{\varepsilon}-U_{*}\|_{\mathrm{op}}\;\leq\;\varepsilon,$ (9)
@@
Proof: The existence of $U_{*}$ with the action (8) follows from Theorem 1.
@@
The two theorems together show that the bit-bit encoding combined with an ansatz that consists of layers of gates from a universal gate set can approximate a classifier to any desired degree of accuracy on a dataset discretized as shown in the previous section. Finally, we observe that there exists a maximum number of bits required to encode the samples in a dataset.
@@ 3.2 Other encoding schemes
Two other encoding schemes are often utilized in the quantum machine learning literature - amplitude encoding, and angle encoding. For both schemes, some kind of compression is required in order to make the data loadable into the quantum computer. Further, within each scheme, the data can be loaded into the quantum computer multiple times in order to enhance the model’s expressivity. However, the combined effects of the compression, data reuploading and design of the variational ansatz on the accuracy that can be attained by the model are not yet understood.
@@
The expressivity of amplitude encoding can be enhanced by loading multiple states in parallel. However, a significant barrier remains in the form of the loading time being proporational to the dimension of the data. Previous work has shown the effectiveness of using tensor network techniques to compress the data before loading it as an amplitude encoded state [26], [4], [27], [28], but has not studied the effect of the compression on the model accuracy.
@@
Angle encoding: Under angle embedding, each input $x\in\mathbb{R}^{n}$ is mapped to a product of single-qubit rotations $U_{\mathrm{enc}}(x)=\bigotimes_{j=1}^{d}R_{\alpha}(\phi_{j}(x))$ (where $R_{\alpha}$ are Pauli rotations), after which a parametrized ansatz $W(\theta)$ is applied and an observable $M$ is measured, yielding $f(x;\theta)=\langle 0|\,U_{\mathrm{enc}}^{\dagger}(x)W^{\dagger}(\theta)MW(\theta)U_{\mathrm{enc}}(x)\,|0\rangle$. Writing $A(\theta)=W^{\dagger}(\theta)MW(\theta)$ and expanding conjugation by $U_{\mathrm{enc}}(x)$ shows that $f$ is a finite trigonometric Fourier series in $x$, namely $f(x;\theta)=\sum_{k}a_{k}(\theta)\cos(\omega_{k}\cdot x)+b_{k}(\theta)\sin(\omega_{k}\cdot x)$ for some frequencies $\omega_{k}\in\mathbb{Z}^{n}$ arising from the Pauli rotation exponents. With $L$ layers of data reuploading, $|\psi(x,\theta)\rangle=\bigl{(}\prod_{\ell=1}^{L}W_{\ell}(\theta_{\ell})U_{\mathrm{enc}}(x)\bigr{)}|0\rangle$, the attainable frequency set expands to $\Omega_{L}\subset\mathbb{Z}^{n}$, so that the expressible function class is the family of trigonometric polynomials $\mathcal{F}_{\mathrm{angle}}^{(L)}=\bigl{\{}\,f(x)=\sum_{\omega\in\Omega_{L}}[a_{\omega}\cos(\omega\cdot x)+b_{\omega}\sin(\omega\cdot x)]:a_{\omega},b_{\omega}\in\mathbb{R}\,\bigr{\}}$.
@@
The set of achievable frequencies $\Omega_{L}$ generated by nested Pauli rotations depends intricately on the circuit architecture and grows in a highly non-regular, combinatorial fashion. Consequently, there is no efficient algorithm that, given a dataset $\mathcal{D}$ or a target function $f^{\star}$, determines the minimum number of reuploading layers $L$ for which $\mathcal{F}_{\mathrm{angle}}^{(L)}$ contains an $\varepsilon$-approximant to $f^{\star}$. Moreover, unlike classical Fourier approximation, where one systematically improves approximation quality by increasing the maximum frequency or degree in a controlled way, the sets $\Omega_{L}$ are not nested in a structured fashion (for example, $\Omega_{L+1}$ does not simply add all frequencies above a certain cutoff), so increasing $L$ does not provide a principled or monotonic method of reducing approximation error. Instead, $L$ must be heuristically chosen, trained, and empirically validated, meaning the universal approximation property does not translate into a systematic, constructive approximation procedure.
@@
In line with this, there are several recent papers that aim to enhance the expressivity of angle encoding. [29] adds auxiliary qubits to enhance the frequency spectrum of an architecture with angle embedding. [30] finds that using data reuploading with circuits that are wider rather than deeper is a more effective strategy. Despite this work on enhancing expressivity, there remains no known method for efficiently and constructively realizing universal approximation with these encoding schemes.
@@
Combining these results, we conclude that bit-bit encoding provides a direct and efficient route to universality in quantum machine learning models. By construction, every dataset of finite precision can be uniquely represented using a bounded number of qubits, and the corresponding classifier can be realized as a unitary operator acting on the joint input–output space. The Solovay–Kitaev theorem then ensures that such unitaries can be approximated to arbitrary accuracy using a finite universal gate set, thereby establishing universality for bit-bit encoded models. In contrast, amplitude and angle encoding require additional structural assumptions or circuit resources in order to achieve comparable universality. Thus, bit-bit encoding yields both a conceptually transparent and practically scalable framework for constructing universally approximating quantum models.
4 Techniques: Bit Encoding Schemes

### v2

@@ Abstract
For a machine learning paradigm to be generally applicable, it should have the property of universal approximation, that is, it should be able to approximate any target function to any desired degree of accuracy. In variational quantum machine learning, the class of functions that can be learned depend on both the data encoding scheme as well as the architecture of the optimizable part of the model. Here, we show that the property of universal approximation is constructively and efficiently realized by the recently proposed bit-bit data encoding scheme. Further, we show that this construction allows us to calculate the number of qubits required to solve a learning problem on a dataset to a target accuracy, giving rise to the first resource estimation framework for variational quantum machine learning. We apply bit-bit encoding to a number of medium-sized classical benchmark datasets and find that they require only $27$ qubits on average for encoding. We extend the basic bit–bit encoding scheme to a variant that efficiently supports batched processing of large datasets. As a demonstration, we apply this new scheme to subsets of a giga-scale transcriptomic dataset. This work establishes bit-bit encoding not only as a universally expressive quantum data representation, but also as a practical foundation for resource estimation and benchmarking in quantum machine learning.
@@ 1 Introduction
The use of quantum computers for machine learning and artificial intelligence applications is a relatively new research area that could provide a means to model datasets which contain correlations that are hard to model classically. Quantum machine learning algorithms have been tested on near-term quantum hardware for various real-world datasets [Gujju24], [elton_copula], [zhu2022copulabased], [iaconis2023tensor], [Johri20], [Rudolph20], [cherrat2022quantum], [silver2023mosaiq], [Silver_Patel_Tiwari_2022]. In particular, variational quantum machine learning, which uses parameterized unitaries that can be trained to model the correlations in a dataset analogous to classical neural networks [Wang_2024], can be applied to a wide variety of problems. However, compared to other applications such as chemistry [Feynman1982], [Lloyd1996],[AspuruGuzik2005],[Reiher2017] or cryptography [Shor1997], where algorithms with well-understood quantum advantage and extensively studied resource estimation are known, there are many more open questions around the use of quantum computers to learn from data.
@@
Among the fundamental questions is one of expressivity — how does one design a variational quantum ansatz, which we refer to as a quantum model, that can approximate the underlying function defined by a given dataset? To answer this question meaningfully, we must note that real-world datasets seldom follow theoretically tractable rules. Therefore, the quantum model should have the property of universal approximation — that is, it should be able to approximate any function with any desired degree of accuracy. Further, the construction which realizes the universal approximation should not itself require exponentially scaling computation, either quantum or classical. If a quantum learning model does not have the property of universal approximation, it is likely not useful in real life. Even if such a model works in a limited setting, we cannot expect it to give good results in the limit of large and complex datasets. Indeed, in classical machine learning, universal approximation theorems [Cybenko1989, Hornik1991] provide the mathematical justification for using neural networks for large and complex datasets.
@@
Schuld et al. [schuld2021effect] showed that the expressivity of a quantum model is constrained by the method used for encoding data into a quantum state. Specifically, a quantum model can be expanded as a partial Fourier series in the encoded data where the frequencies present in the series are set by linear combinations of the eigenvalues of the encoding operators. Following this, it is possible to increase the number of frequencies in the series by reusing the data multiple times in the quantum ansatz, a technique known as data re-uploading [perez2020data]. The time for loading is then lower-bounded by the quantity of information to be loaded multiplied by the number of times it is loaded. With this method, even an image from a simple dataset like MNIST, which has 784 pixels, is practically expensive to load directly into a quantum state. Since this limit is fundamental, it is clear that some kind of compression is required before loading, for which techniques like principal component analysis are often used. However, it remains unclear how much the degree of compression and the number of data re-uploadings affect the accuracy that can be reached by the model.
@@
In this work, we show that the bit-bit encoding technique [johri2025bitbitencoding], which was recently proposed for supervised quantum learning, directly addresses the problem of achieving efficient compression with predictable accuracy. In this encoding, both the input and the output are encoded as binary strings following a dimensionality reduction procedure. These binary strings correspond to the basis states of a quantum computer at initialization and measurement. For a small number of qubits, there are collisions in which the same input corresponds to more than one output, but as the number of qubits increases, the encoding becomes more precise. In this limit, we prove that a variational quantum model with bit-bit encoding has the property of universal approximation. Further, the classical pre-processing required before data loading utilizes techniques such as principal component analysis which scale only polynomially in the dataset size.
@@
A useful emergent property of the bit-bit encoding construction is that it is possible to efficiently calculate the number of qubits required to encode a dataset to a desired degree of accuracy. This then gives us the first framework for resource estimation in quantum machine learning. Using this framework, we can, for the first time, estimate the number of qubits required to successfully train a model on a given dataset using a variational approach. In turn, this analysis lets us assess the potential for quantum advantage on a dataset as we can calculate whether a learning problem fits into a number of qubits that is classically simulatable.
@@
We introduce $Q_{\text{dataset}}(x)$, a resource metric that estimates the number of qubits required to model a dataset to accuracy $x$ using bit-bit encoding; $Q_{\text{dataset}}(1.0)$ is the number of qubits required to achieve both $100\%$ training and testing accuracy. Starting with the MNIST dataset, we illustrate how $Q_{\text{dataset}}(x)$ is derived by incrementally increasing qubit allocations until collisions on both the train and test datasets vanish, revealing a trade-off between maximizing theoretical training accuracy and achieving generalization. We find that medium-sized classical benchmark classification datasets have a $Q_{\text{dataset}}(1.0)$ of $27$ on average. Comparisons across dimensionality reduction schemes show that bit-bit encoding achieves more efficient data compression with schemes that produce independent or de-correlated features. Finally, scaling experiments on a transcriptomic dataset reveal examples where subsets of the dataset have a much higher $Q_{\text{dataset}}(1.0)$ than the benchmark datasets despite having comparable dimensionality, indicating that large transcriptomic dataset learning problems are promising candidates for quantum advantage.
@@
We note that we only consider classical datasets in this work. Although there has been progress in quantum machine learning for quantum data [Nagano2023QuantumDataLearning], [Huang2022QuantumAdvantageExperiments], practical use cases remains limited. In contrast, quantum machine learning applied to classical data offers a far broader range of potential applications.
@@
The paper is organized as follows. We begin with an overview of bit-bit encoding in Section 2, followed by an analysis of expressive power and universal approximation in commonly used encoding schemes in Section 3. Section 4 presents calculations of $Q_{\text{dataset}}$ on classical benchmark datasets across dimensionality reduction schemes along with an extension of bit-bit encoding for large-scale data encoding. Section 5 demonstrates training of a classifier on a bit-bit encoded dataset. We conclude with a discussion in Section 6.
2 Methods
2.1 Overview of Bit-Bit Encoding
@@
Bit-bit encoding [johri2025bitbitencoding] encodes datasets with numeric features and discrete class labels into bitstrings, which allows the dataset to be loaded into the Hilbert space of a given number of qubits for classification. Consider a supervised learning problem for a dataset with a feature matrix $X\in\mathbb{R}^{s\times n}$ and class label vector $y\in\{1,2,\ldots,c\}^{s}$, where $s$ is the number of samples, $n$ is the number of features, and $c$ is the number of class labels. For a classification problem, the learning task consists of finding an efficiently computable function that maps $X$ to $y$. Typically, the function is many-to-one, that is, $c<|X|$.
@@
Bit-bit encoding converts the feature matrix $X$ into a discretized feature matrix $Z$, compactly encoded into $N_{x}$ qubits. Bit-bit encoding works as follows. First, the dataset is split into train and test sets: $X_{\text{train}}$, $y_{\text{train}}$, $X_{\text{test}}$, and $y_{\text{test}}$. A dimensionality reduction model, principal component analysis (PCA) by default, is fit on $X_{\text{train}}$ and applied to $X_{\text{test}}$, reducing both sets to $D$ components. The number of bits assigned to each component is calculated with
@@
where $I$ is the mutual information and $N_{x}$ is the number of bits allocated to the discretized feature matrix. The dimension-reduced $X_{\text{train}}$ and $X_{\text{test}}$ are then min-max normalized so that each sample lies in $0,1]^{D}$. A copula transformation is fit on the reduced and normalized $X_{\text{train}}$ and applied to the reduced and normalized $X_{\text{test}}$, mapping each component to its empirical cumulative distribution function so that the feature distributions are approximately uniform. Finally, to encode the reduced, normalized, and transformed $X_{\text{train}}$ and $X_{\text{test}}$, each data point $X_{sd}$ (the $s^{th}$ sample of the $d^{th}$ feature) is discretized into $b_{d}$ bits of precision with
@@
Discretizing the feature matrix also compresses it, potentially creating collisions at the input as a particular feature vector $z\in Z$ may occur more than once, each time corresponding to the same or different value of $y_{s}\in y$. The samples in the discretized feature matrix can then be considered as samples from a joint probability distribution $f(Z=z,y=y_{s})$ over correlated random variables $(Z,y)$. For the purpose of classification, the “correct” output is then taken to be the mapping which occurs most frequently, that is,
@@
Given this form of the classification function, the quantum learning problem can be cast in terms of an operator $U_{*}$ that has the following action on $N_{q}=N_{x}+N_{y}$ qubits:
@@
A linear loss function can be constructed as
$\bar{L}(\vec{\theta})=1-\sum_{z\in\mathcal{D}_{\vec{b}}}f(z)P_{C(z),z}(\vec{\theta}),$ (2.3)
where $0\leq\bar{L}\leq 1$ and $f(z)$ is the frequency of occurrence of $z$ in the dataset. This loss function returns the probability of the model giving the incorrect answer when queried. Similarly, non-linear loss functions can also be constructed. If $U(\vec{\theta})=U_{*}$ and there is no noise, even one shot would suffice to classify the input data sample.
We distinguish the bit-bit encoding technique from a simple basis encoding as it can be used to represent datasets that do not originally have a binary representation, but are converted into binary representation through a process of dimensionality reduction followed by mutual-information based discretization. Indeed, a basis encoding that utilized only a naive discretization of the dataset would require many more qubits. For instance, basis encoding a dataset consisting of molecular fingerprint samples [boy2025encodingmolecularstructuresquantum], which are naturally binary, would take thousands of qubits.
@@ 3.1 Universality of bit-bit encoding
In this section, we show how the bit-bit encoding leads to a universal approximation theorem for classification. First, we prove that the classification operator $U_{*}$ defined in Eq. 2.1 is unitary.
@@
$C(z)=C(z^{\prime})\;\Longrightarrow\;\braket{g(z)|g(z^{\prime})}=\delta_{z,z^{\prime}}.$ (3.1)
@@
If $C(z)\neq C(z^{\prime})$ the first factor vanishes; if $C(z)=C(z^{\prime})$, the condition (3.1) gives $\braket{g(z)|g(z^{\prime})}=\delta_{z,z^{\prime}}$. Hence, in all cases,
$\bigl\langle C(z),g(z)\bigm|C(z^{\prime}),g(z^{\prime})\bigr\rangle=\delta_{z,z^{\prime}}.$ (3.4)
@@
is an orthonormal family in $\mathcal{H}$. Because $|\{0,1\}^{N_{x}}|=2^{N_{x}}$ and $\dim\mathcal{H}_{\mathrm{in}}=2^{N_{x}}$, we have $\dim\mathrm{span}(\mathcal{B}_{\mathrm{out}})=2^{N_{x}}$. Thus $V$ in (3.2) is an isometry from $\mathcal{H}_{\mathrm{in}}$ onto
@@
$\|U_{\varepsilon}-U_{*}\|_{\mathrm{op}}\;\leq\;\varepsilon,$ (3.6)
@@
Proof: The existence of $U_{*}$ with the action (3.5) follows from Theorem 1.
@@
The two theorems together show that the bit-bit encoding combined with an ansatz that consists of layers of gates from a universal gate set can approximate a classifier to any desired degree of accuracy on a dataset discretized as shown in the previous section.
Finally, we observe that there exists a maximum number of bits required to encode the samples in a dataset.
@@ 3.2 Other encoding schemes
Two other encoding schemes are often utilized in the quantum machine learning literature — amplitude encoding, and angle encoding. For both schemes, some kind of compression is required in order to make the data practically loadable into the quantum computer. Further, within each scheme, the data can be loaded into the quantum computer multiple times in order to enhance the model’s expressivity. However, there does not exist a framework to understand the combined effects of the compression, data re-uploading and design of the variational ansatz on the approximation power of the model.
@@
The expressivity of amplitude encoding can be enhanced by loading multiple states in parallel. However, a significant barrier remains in the form of the loading time being proporational to the dimension of the data. Previous work has shown the effectiveness of using tensor network techniques to compress the data before loading it as an amplitude encoded state [iaconis2023quantum, iaconis2023tensor, pollmann_tensor, jumade2023dataloadableshortdepth], but has not studied the effect of the compression on the model accuracy.
@@
Angle encoding: Under angle embedding, each input $x\in\mathbb{R}^{n}$ is mapped to a product of single-qubit rotations $U_{\mathrm{enc}}(x)=\bigotimes_{j=1}^{d}R_{\alpha}(\phi_{j}(x))$ (where $R_{\alpha}$ are Pauli rotations), after which a parametrized ansatz $W(\theta)$ is applied and an observable $M$ is measured, yielding $f(x;\theta)=\langle 0|\,U_{\mathrm{enc}}^{\dagger}(x)W^{\dagger}(\theta)MW(\theta)U_{\mathrm{enc}}(x)\,|0\rangle$. Writing $A(\theta)=W^{\dagger}(\theta)MW(\theta)$ and expanding conjugation by $U_{\mathrm{enc}}(x)$ shows that $f$ is a finite trigonometric Fourier series in $x$, namely $f(x;\theta)=\sum_{k}a_{k}(\theta)\cos(\omega_{k}\cdot x)+b_{k}(\theta)\sin(\omega_{k}\cdot x)$ for some frequencies $\omega_{k}\in\mathbb{Z}^{n}$ arising from the Pauli rotation exponents. With $L$ layers of data re-uploading, $|\psi(x,\theta)\rangle=\bigl(\prod_{\ell=1}^{L}W_{\ell}(\theta_{\ell})U_{\mathrm{enc}}(x)\bigr)|0\rangle$, the attainable frequency set expands to $\Omega_{L}\subset\mathbb{Z}^{n}$, so that the expressible function class is the family of trigonometric polynomials $\mathcal{F}_{\mathrm{angle}}^{(L)}=\bigl\{\,f(x)=\sum_{\omega\in\Omega_{L}}[a_{\omega}\cos(\omega\cdot x)+b_{\omega}\sin(\omega\cdot x)]:a_{\omega},b_{\omega}\in\mathbb{R}\,\bigr\}$.
@@
The set of achievable frequencies $\Omega_{L}$ generated by nested Pauli rotations depends intricately on the circuit architecture and grows in a highly non-regular, combinatorial fashion. Consequently, there is no efficient algorithm that, given a dataset $\mathcal{D}$ or a target function $f^{\star}$, determines the minimum number of re-uploading layers $L$ for which $\mathcal{F}_{\mathrm{angle}}^{(L)}$ contains an $\varepsilon$-approximant to $f^{\star}$. Moreover, unlike classical Fourier approximation, where one systematically improves approximation quality by increasing the maximum frequency or degree in a controlled way, the sets $\Omega_{L}$ are not nested in a structured fashion (for example, $\Omega_{L+1}$ does not simply add all frequencies above a certain cutoff), so increasing $L$ does not provide a principled or monotonic method of reducing approximation error. Instead, $L$ must be heuristically chosen, trained, and empirically validated, meaning the universal approximation property does not translate into a systematic, constructive approximation procedure.
@@
In line with this, there are several recent papers that aim to enhance the expressivity of angle encoding. [wen2024enhancingexpressivityquantumneural] adds auxiliary qubits to enhance the frequency spectrum of an architecture with angle embedding. [wang2025predictiveperformancedeepquantum] finds that using data re-uploading with circuits that are wider rather than deeper is a more effective strategy. Despite this work on enhancing expressivity, there remains no known method for efficiently and constructively realizing universal approximation with these encoding schemes.
@@
Combining these results, we conclude that bit-bit encoding provides a direct and efficient route to universality in quantum machine learning models. By construction, every dataset can be represented using a bounded number of qubits, and the corresponding classifier can be realized as a unitary operator acting on the joint input–output space. The Solovay–Kitaev theorem then ensures that such unitaries can be approximated to arbitrary accuracy using a finite universal gate set, thereby establishing universality for bit-bit encoded models. In contrast, there is no practical known construction for amplitude and angle encoding to achieve comparable universality. Thus, bit-bit encoding yields both a conceptually transparent and practically scalable framework for constructing universally approximating quantum models.
4 Techniques & Results: Bit-bit Encoding Schemes